### v2

# Surfel-based Gaussian Inverse Rendering for Fast and Relightable Dynamic Human Reconstruction from Monocular Videos

Abstract
Efficient and accurate reconstruction of a relightable, dynamic clothed human avatar from a monocular video is crucial for the entertainment industry. This paper introduces the Surfel-based Gaussian Inverse Avatar (SGIA) method, which introduces efficient training and rendering for relightable dynamic human reconstruction. SGIA advances previous Gaussian Avatar methods by comprehensively modeling Physically-Based Rendering (PBR) properties for clothed human avatars, allowing for the manipulation of avatars into novel poses under diverse lighting conditions. Specifically, our approach integrates pre-integration and image-based lighting for fast light calculations that surpass the performance of existing implicit-based techniques. To address challenges related to material lighting disentanglement and accurate geometry reconstruction, we propose an innovative occlusion approximation strategy and a progressive training approach. Extensive experiments demonstrate that SGIA not only achieves highly accurate physical properties but also significantly enhances the realistic relighting of dynamic human avatars, providing a substantial speed advantage. We exhibit more results in our project page: https://GS-IA.github.io.
Index Terms:
Relightable Dynamic Human Reconstruction, PBR Properties Estimation, Novel View Synthesis
1 Introduction
Figure 1: We achieve fast reconstruction of clothed human avatars with PBR properties from a monocular video. SGIA takes a monocular video and initial human pose and shape as input to estimate dynamic clothed humans’ PBR properties, including geometry and materials. Leveraging a PBR-aware 2DGS representation, our method enables fast training and rendering processes. By utilizing the estimated PBR properties, we can not only deform the avatars into different poses but also render them with realistic lighting conditions, allowing for versatile and visually appealing outputs.
Reconstructing a clothed human avatar and estimating its Physically-Based Rendering (referred to as PBR) properties from monocular videos is a challenging task in computer vision and graphics. This process involves disentangling geometry, PBR materials, and environment illumination. With accurate PBR properties, avatars can be deformed into different poses and re-rendered under novel illuminations, as shown in Fig. 1, offering significant potential for virtual reality, movies, and games.
Early works aimed to estimate PBR properties from single images [1, 2, 3, 4, 5, 6, 7] or multi-view inputs [8, 9, 10, 11, 12, 13, 14, 15], often assuming known lighting environment maps or scene geometry, which is impractical. Radiance fields [16] have emerged as effective 3D-consistent scene representations. Works in this direction [17, 18, 19, 20, 21, 22, 23, 24] estimate geometry, material, and environment light for static scenes from multi-view images but do not account for dynamics, crucial for modeling clothed human avatars. Recent research [25, 26, 27] has advanced clothed human animation by modeling radiance fields and PBR properties for dynamic humans, but NeRF optimizes density rather than geometry directly [28], leading to suboptimal geometry quality. Some works explore neural implicit surfaces [29, 30] for better geometry, but all these methods rely on dense point sampling from MLPs [16] or multi-resolution hash-grids [31], slowing down training and rendering due to extensive query operations.
Recently, Gaussian Splatting [32, 33] has revolutionized high-quality, real-time rendering through its explicit representation and efficient splatting-based rasterization on modern GPUs. Unlike prior methods [34, 35, 29] that depend on complex backward skinning [36, 37] for transformations to a canonical space [38], Gaussian primitives for clothed human avatars are easily and efficiently animated via forward skinning, similar to the template mesh [39]. Recent studies [40, 41, 42, 43, 44, 45, 46] have applied Gaussian splatting techniques to reconstruct dynamic clothed human avatars from monocular or multi-view videos, achieving remarkable rendering quality and speed. However, these approaches are limited to capturing radiance fields and fail to recover the full Physically-Based Rendering (PBR) properties of dynamic humans, thus lacking generalization to novel illuminations.
Our goal is to efficiently and accurately reconstruct a relightable, dynamic clothed human avatar from a monocular video, addressing two key challenges: i.e., disentangling coupled materials and shadow effects caused by lighting occlusion, and achieving high-quality geometry of the clothed human avatar. To this end, we propose the Surfel-based Gaussian Inverse Avatar (SGIA), inspired by the efficiency of Gaussian representation. SGIA recovers both the radiance field and Physically-Based Rendering (PBR) properties of clothed human avatars from monocular videos, enabling avatars to adapt to any scene and varying lighting conditions. We define a set of canonical PBR-awre 2DGS in a canonical space, which are transformed into the observation space via forward linear blend skinning based on different poses. To accurately model the dynamics of clothed human avatars, we combine articulation from the template model Skinned Multi-Person Linear Model (SMPL)[39] with learnable latent bones conditioned on human poses, drawing inspiration from[40].
To effectively disentangle the coupled materials and shadow effects caused by lighting occlusion, inspired by previous works [47, 48, 49, 50], we add learnable parameters, such as albedo, roughness, and metallic, to each Gaussian primitive for the clothed human avatar. While occlusion modeling can be efficiently handled with spherical harmonics in static scenes, it becomes challenging in dynamic scenes where occlusion relationships change over time. Directly calculating occlusion with ray tracing [51] for each Gaussian is accurate but inefficient. To achieve efficient occlusion calculations, we use modern ray-casting algorithms on the limited vertices of the template mesh. Our hybrid strategy combines point and mesh benefits: querying the nearest points on the template mesh for each Gaussian and using surface points and normal vectors to bake ambient occlusion from the template mesh. Given the minimal occlusion difference between the clothed human and the unclothed SMPL [39], our method balances efficiency and accuracy, effectively disentangling materials and lighting.
To accurately reconstruct the surface geometry of clothed human avatars, which is crucial for physically based inverse rendering, we propose to leverage the 2D Gaussian Splatting (referred to as 2DGS) [33] as our primitive representation due to its surface-aligned nature, improving geometry reconstruction quality. Each 2D surfel’s orientation serves as a local normal direction. Initially, we optimized the rough shape using the original Gaussian color. However, the geometry remains inaccurate after image reconstruction, especially under high light conditions (Fig.3). Unlike previous works[47, 50, 49] that freeze Gaussian shape after the first stage, we refine the shape attributes during the PBR optimization stage, leveraging the normal-correlated appearance to enhance geometry. While the original normal consistency loss in 2DGS maintains a splat normal consistency with geometry, it can hinder normal optimization. To address this, we propose a progressive optimization strategy that achieves both accurate geometry and consistent splat normals.
By utilizing the materials, occlusion, and geometry for each splat, we calculate the PBR color of each Gaussian in the observation space and apply the splatting process [52, 33] designed for 2DGS to produce the PBR image. To address the ambiguity between lighting and materials and achieve plausible PBR material solutions, we use regularization terms to maintain material smoothness on planar surfaces and ensure natural lighting. Extensive experiments on synthetic and real-world datasets show that our method enables fast, relightable clothed human avatar reconstruction within 40 minutes and renders at $540\times 540$ resolution at 5 FPS, which is 100×100$\times$100 × faster than state-of-the-art NeRF-based human inverse rendering methods. Quantitative evaluations demonstrate that our method achieves comparable PBR property estimation accuracy to state-of-the-art methods. Comprehensive ablation studies further validate our approach.
In summary, our main contributions are:
  • We propose SGIA, a surfel-based PBR properties modeling method for fast and relightable dynamic clothed human reconstruction from monocular videos.
  • Our method tackles the challenges in material lighting disentanglement and accurate geometry reconstruction by utilizing our proposed occlusion approximation strategy and progressive training strategy.
  • Extensive experiments demonstrate that our method achieves fast training and rendering while maintaining high-accuracy in PBR properties estimation and human relighting at novel poses.
2 Related Work
2.1 Inverse rendering for static scenes.
Estimating the physical properties from images is a highly indeterminate problem due to its under-constrained nature. Previous works on BTF [53, 54, 55, 56, 57] and SVBRDF [58, 11, 14, 59] estimation relay on specific assumptions of specific viewpoints, lighting patterns, or capture setups. Other works explore inverse rendering on single images [1, 2, 3, 4, 5, 6, 7] or multi-view inputs [8, 9, 10, 11, 12, 13, 14, 15]. Recovering shape and illumination from RGB images is heavily ill-posed due to the coupling of shape, materials, and lighting in real-world captured images. These works can only recover high-quality material for each view but fail to achieve realistic and consistent multi-view relighting results under real-world natural illuminations due to the lack of 3D consistent material fields.
With the rapid development of Neural Radiance Fields (referred to as NeRFs) [16], lots of works [23, 17, 18, 19, 21, 20, 60, 61, 22, 62, 63, 64] based on neural field have been proposed to achieve inverse rendering from multi-view images. Specifically, NeRFactor [23] enabled full estimation of the physical properties (geometry, albedo, materials, and lighting) by incorporating the effect of shadows. Neural-PIL [21] explored to approximate pre-integration light integration process by a neural network [65] to achieve fast calculation of outgoing radiance. PhySG [22] proposed to use mixtures of spherical Gaussians to approximate the light transport efficiently. InvRender [62] proposed to model the indirect illumination by distilling a pre-trained NeRF into MLPs to achieve more accurate light disentangling. TensoIR [24] took advantage of tensor decompostion [66] to improve the reconstruction quality. Due to the inaccurate geometry modeling in NeRFs, some other works [67, 68, 69, 70, 71] explored neural signed distance function (SDFs) rendering to achieve more accurate geometry reconstruction during inverse rendering. Among them, NeRO [67] applied the split-sum to constrain the learning of reflective objects, and it can achieve high-quality geometry reconstruction for the high reflective objects. The deep Marching Tetrahedra [72] provided an efficient and differentiable mesh extraction from neural implicit surface representation. With the benefit of differentiable mesh representation and the differentiable mesh rasterizer, Nvdiffrec [73] proposed to conduct light integration on the mesh surface to achieve rapid and photo-realistic physically based rendering. NvdiffrecMC [74] showed that by exploring the combination of mesh-based Mote-Carlo ray-tracing and multi-importance sampling, the decomposition into shape, materials, and lighting can be further improved.
More recently, Gaussian Splatting based methods [32, 75, 33, 76] have achieved high-quality rendering and real-time rendering speed thanks to the forward mapping splatting process. Some works [47, 48, 49, 50] explored to use either the learnable normal vector [47, 50] or the shortest axis of the Gaussian [48, 49] to achieve geometry representation for inverse rendering. However, these works only focused on the inverse rendering of the static scene, while our work focuses on the inverse rendering of a dynamic clothed human avatar, which is more challenging due to the avatar’s dynamic.
2.2 Clothed human avatar modeling.
In addition to the textured mesh, NeRFs [16] also became a useful representation for photo-realistic clothed avatar reconstruction [77, 38, 78, 79, 80, 35, 34, 81, 82, 83, 84] from monocular or multi-view videos. NeuralBody [38] assigned the latent code on the canonical space SMPL [39]vertices and applied the LBS [39] to transform the clothed human avatars to articulation space. For the ray sampling in NeRF [16], the world space points are first transformed to articulated SMPL space to query the radiance. They also rely on time-varied latent to fit the pose-dependent deformation. Animatable NeRF [77] further proposed a pose-dependent deformation network to fit the clothed human conditioned on human. With pose-dependent modeling, the reconstructed avatar can be generalized to novel poses. They also utilize inverse LBS, which can transform the sample points from SMPL articulated space to the canonical space to query the radiance, thus achieving accurate and fast radiance field modeling for humans. However, there exists a large ambiguity in the inverse LBS calculation. InstantAvatar [35] proposed to incorporate the accurate and root finding [37, 36] for inverse LBS and multi-resolution Hash Grids for radiance fields modeling. Recent advances in Gaussian Splatting [32] assigned the radiance to explicit points that can be easily animated by manipulating the skinning weight from template mesh. The forward splatting process does not need the transformation from articulated space to canonical space. Therefore, both the ambiguous inverse LBS and time-consuming root-finding in Snarf can be avoided. Most works [43, 40, 44, 85, 86] explored similar techniques to achieve fast clothed human avatar reconstruction from monocular videos. In contrast with these works, our work aims to model the PBR properties rather than only the radiance. We also take the disentangling of geometry, materials, and environmental lighting to achieve estimation of full PBR properties.
2.3 PBR properties reconstruction of clothed human avatars.
Previous works [87, 88, 89, 90] explored neural networks to learn the decomposition of materials and lighting from single portrait images. Recently, High quality physical properties reconstruction of clothed human avatar [30, 91, 25, 26, 29, 27, 92, 93, 94, 79] have been widely explored with the development of Neural Radiance Fields [16]. DS-NeRF [79] proposed to query the lighting at world space to disentangle the light and texture. Relighing4D [25] first explored the MLPs representation based on NeuralBody [38] to jointly estimate the shape, lighting, and the albedo of dynamic humans from monocular or sparse videos under unknown illuminations by disentangling visibility via MLPs. RANA [92] pre-trained a mesh representation on ground truth 3d assets to approximate the radiance transfer under different lighting. Relightable Avatar [26] pretrained an occlusion approximation neural network from a large-scale motion sequence dataset [95, 96]. Intrinsic Avatar [29] implemented a fast secondary ray-tracing at articulated space to calculate the visibility and the indirect illuminations. Different from these works, our method explores a relightable clothed human avatar representation based on Gaussian splatting [32] to achieve fast training and fast rendering. Concurrent with our work, Animatable Gaussians [97] proposed a relighting extension that predicts occlusion based on the human pose and position map with the 2D StyleUnet [98] to achieve high-quality rendering. However, their work relied on multi-view videos as input and complex poses during training to achieve pose generalization, which is not suitable for our setting. PhysAvatar [99] also explored the relightable Gaussian Avatar, but it relied on the multi-view videos as input to reconstruct the mesh at the first timestamp.
3 Our Method
Figure 2: Overview of our pipeline. We define the radiance ($c_{k}$) and materials ($a_{k},r_{k},m_{k}$) at the canonical space as canonical PBR-aware 2DGS and deform them to the world space via Linear Blend Skinning (LBS). We first optimize with Image reconstruction loss to get the initial shape of the clothed avatar. Based on the rough shape, we optimize the Gaussian attributes with Physically Based Rendering. To model the shadow effect and decouple the materials from lighting, we propose to approximate the calculation of occlusion from the template mesh. Additionally, we apply regularization, i.e., smooth loss, white loss, and normal consistency loss to get a plausible solution for the PBR materials.
In this section, we first introduce the background of surfel-based representation [33, 32]. Then, we describe our framework for relightable dynamic clothed avatar representation and physically-based rendering. The clothed avatar is built upon surfel-based representation [33] with an articulated template [40] based on SMPL [39]. Motivated by the success of 2DGS for high-quality geometry recovery, we propose a PBR-aware 2DGS that incorporates the lighting, roughness, metallics, and albedo into the 2DGS for monocular videos. We also propose a solution for ambient occlusion calculation to enhance the rendering realism and better decoupling of the material and lighting. Finally, we describe our progressive training strategy to achieve geometrically accurate surface reconstruction and accurate albedo estimation. We visualize the overall pipeline of our method in Fig. 2.
3.1 A revisit of 2DGS
Unlike 3DGS [32], 2DGS [33] adopt flat 2D Gaussians embedded in 3D space for scene representation. The primitive distributes densities within a planar disk (also known as surfel), which makes the primitive better alignment with surfaces. This feature can improve the quality of our geometry reconstruction and achieve better material estimation. In 2DGS [33], each primitive is characterized by a central points $\mathbf{p}_{k}\in\mathbb{R}^{3}$ , two principal tangential vectors $\mathbf{t}_{u},\mathbf{t}_{v}\in\mathbb{R}^{3}$ and two scaling factors $s_{u},s_{v}\in\mathbb{R}$. The scaling factors correspond to the variances of the 2D Gaussian.
A 2D Gaussian is defined in a local tangent plane of the world space, parameterized as :
$\mathbf{P}(u,v)=\mathbf{p}_{k}+s_{u}\mathbf{t}_{u}u+s_{v}\mathbf{t}_{v}v=[\mathbf{R}_{k}\mathbf{S}_{k},\mathbf{p}_{k}](u,v,1,1)^{T}$ (1)
where $\mathbf{R}_{k}=[\mathbf{t}_{u},\mathbf{t}_{v},\mathbf{t}_{u}\times\mathbf{t}_{v}]\in SO(3)$ denotes the rotation of each surfel, $\mathbf{t}_{u}\times\mathbf{t}_{v}$ denotes the normal orientation of each surfel, and the diagonal matrix $\mathbf{S}_{k}=\text{diag}(s_{u},s_{v},0)\in\mathbb{R}^{3\times 3}$ denotes the scale of each surfel. Each 2D Gaussian primitive also carries an opacity $o$ and a view-dependent appearance $\mathbf{c}$.
For each point $\mathbf{u}=(u,v)$ in the local coordinate, its 2D Gaussian value can be calculated by
$\mathcal{G}(\mathbf{u})=\exp(-\frac{u^{2}+v^{2}}{2})$ (2)
Then, 2D Gaussians are sorted based on the depth of their center and organized into tiles. To render a pixel $x$, we then cast a ray into the scene to compute the intersection $\mathbf{u}(x)$ on the splats and accumulate the texture function with volumetric alpha blending:
$\mathbf{c}(x)=\sum_{i=1}\mathbf{c}_{i}o_{i}\mathcal{G}_{i}(\mathbf{u}(x))\prod_{j=1}^{i-1}(1-o_{j}\mathcal{G}_{j}(\mathbf{u}(x)))$ (3)
3.2 Clothed humans avatars as dynamic surfels animated by template models
We follow the recent advances [40, 100, 44, 41] to model the human as animatable Gaussians. Specifically, we define the PBR-aware 2DGS in a canonical space and transform them into world space via linear blend skinning (LBS). The Gaussians are animated by a human template model [39] to achieve avatar animation. In contrast to them, we choose the surfel-based primitive [33] as our representation for better geometry accuracy. Furthermore, we define additional parameters for intrinsic properties and lighting on each 2D Gaussian to enable physically based rendering.
Canonical PBR-aware 2DGS Representations. We first utilize a set of PBR-aware 2D Gaussians ($\mathbf{G}$) with material parameters to represent the clothed avatar in a canonical space:
$\mathbf{G}=\{\mathbf{p}_{k},\mathbf{R}_{k},\mathbf{S}_{k},\mathbf{c}_{k},o_{k},\mathbf{a}_{k},r_{k},m_{k}\}_{k=1}^{N_{\text{gs}}}$ (4)
$N_{\text{gs}}$ is the number of the gaussians. $\mathbf{p}_{k},\mathbf{R}_{k},\mathbf{S}_{k},o_{k}$ is the parameters for the shape of each Gaussian. In the first stage, we utilize $\mathbf{c}_{k}$ as appearance to optimize the shape parameters of the Gaussians. At the second stage, to achieve PBR materials estimation and physically based rendering, we additionally introduce parameters albedo $\mathbf{a}_{k}\in\mathbb{R}^{3}$, roughness $r_{k}\in\mathbb{R}$ and metallics $m_{k}\in\mathbb{R}$ to model the decoupled appearance on each 2D Gaussian. We utilize $N_{\text{init}}=6980$ canonical SMPL template vertices and their corresponding normal vectors to initialize the location and rotation of our canonical Gaussians. The initialization of scaling factors is calculated in the canonical space following [40].
Articulation via LBS. We utilize SMPL [39] as our articulation template model. The SMPL model is defined by a function $M(\boldsymbol{\theta},\boldsymbol{\beta}):\mathbb{R}^{|\boldsymbol{\theta}|\times|\boldsymbol{\beta}|}\rightarrow\mathbb{R}^{3N_{\text{init}}}$, where the $\boldsymbol{\theta}\in\mathbb{R}^{3\times K}$ ($K$ denotes the number of body joints) denotes the body pose parameters, $\boldsymbol{\beta}\in\mathbb{R}^{10}$ denotes the coefficients for body shape.
With the template model, the Gaussian points can be deformed to the articulated space via linear blend skinning (LBS). For each canonical space Gaussian point $\mathbf{p}_{k}$, we first calculate the articulation transformation $\mathbf{A}_{k}=[\mathbf{A}_{k}^{rot},\mathbf{A}_{k}^{t}]$ based on the pose $\boldsymbol{\theta}$ where $\mathbf{A}_{k}^{rot}$ denotes the rotation matrix and $\mathbf{A}_{k}^{t}$ denotes the translation vector for the articulated transformation. The $\mathbf{A}_{k}$ for each Gaussian can be calculated via the following equation:
$\mathbf{A}_{k}=\sum_{i=1}^{N_{b}}\widehat{\mathbf{W}}_{i}(\mathbf{p}_{k})\mathbf{B}_{i}$ (5)
$\mathcal{B}(\boldsymbol{\theta})=[\mathbf{B}_{1},\mathbf{B}_{2},\dots,\mathbf{B}_{N_{b}}]$ ($\mathbf{B}_{i}\in SE(3)$) denotes the rigid transformation that moves each canonical joint to the articulated space. $N_{b}$ is the number of joints in a template model. $\widehat{\mathbf{W}}(\mathbf{p}_{k})$ denotes the blending weight for the bones $\mathcal{B}(\boldsymbol{\theta})$. We model with $\widehat{\mathbf{W}}(\mathbf{p}_{k})=\mathbf{W}(\mathbf{p}_{k})+\Delta\mathbf{W}(\mathbf{p}_{k})$. $\mathbf{W}(\mathbf{p}_{k})\in R^{N_{b}}$ is the predefined skinning weight queried in canonical space for each Gaussian, and the $\Delta\mathbf{W}(\mathbf{p}_{k})$ is the learnable blending weight offset for each Gaussian. We follow [40] and add this term to the predefined blending $\mathbf{W}$ to model the actual instance deformation. We follow Instant Avatar [35] to store the predefined skinning weight of canonical space in a dense 3D volume. Then, we can deform the 2D Gaussians to the articulated space via:
$\mathbf{p}_{k}^{*}=\mathbf{A}_{k}^{rot}\mathbf{p}_{k}+\mathbf{A}_{k}^{t},\quad\mathbf{R}_{k}^{*}=\mathbf{A}_{k}^{rot}\mathbf{R}_{k}.$ (6)
where $\mathbf{p}_{k}^{*}$ and $\mathbf{R}_{k}^{*}$ denotes the location and rotation of 2D Gaussians in articulated space.
To approximate the clothing motion that can be captured from the monocular video, we follow [40] and introduce $N_{\text{lb}}=4$ latent Bones ~$\widetilde{\mathcal{B}}(\boldsymbol{\theta})$ parameterized with an MLP conditioned on human pose $\boldsymbol{\theta}$. We also parameterized the blending weight $\widetilde{\mathbf{W}}_{i}$ for each latent bones ~$\widetilde{\mathcal{B}}_{k}(\boldsymbol{\theta})$. We therefore extend the equation 5 to the overall articulation transformation:
$\mathbf{A}_{k}=\sum_{i=1}^{N_{b}}\widehat{\mathbf{W}}_{i}(\mathbf{p}_{k})\mathbf{B}_{i}+\sum_{i=1}^{N_{lb}}\widetilde{\mathbf{W}}_{i}(\mathbf{p}_{k})\widetilde{\mathbf{B}}_{i}$ (7)
3.3 Physically based rendering with image-based Lighting
Physically Based Inverse Rendering. Unlike previous physically-based cloth avatars [29, 30], our lighting calculation does not need the time-consuming Monte Carlo Sampling, the integration of ingoing radiance in the hemisphere. We instead utilize the image-based lighting and split-sum approximation [101] to tackle the intractable integral.
The rendering equation involves the calculation of the outgoing light from a surfel at observation space $\mathbf{p}_{k}^{*}$  [^1] We omit $\mathbf{p}_{k}$ for simplified representation. All the $\mathbf{L}_{o},\mathbf{L}_{i},\mathbf{n},\mathbf{a},m$ in the following equation denotes the corresponding parameters on each surfel $\mathbf{p}^{*}_{k}$ based on the BRDF materials, which can be written as:
[…]
where $\mathbf{w}_{o}$ denotes the view direction of outgoing radiance $\mathbf{L}_{o}(\mathbf{w}_{o})$, $\mathbf{w}_{i}$ denotes the direction of the ingoing radiance $\mathbf{L}_{i}(\mathbf{w}_{i})$ sampled from the hemisphere $\Omega$, $\mathbf{n}$ denotes the normal orientation of the surfel. The BRDFs property $f$ can be divided into a specular term (the 1st term in equation. 9) and a diffuse term (the 2nd term in equation. 9) following the Disney-BRDF model [102]:
$f(\mathbf{w}_{o},\mathbf{w}_{i})=\frac{DFG}{4(\mathbf{w}_{i}\cdot\mathbf{n})(\mathbf{w}_{o}\cdot\mathbf{n})}+(1-m)\frac{\mathbf{a}}{\pi}$ (9)
where $D$ denotes the GGX [103] normal distribution function (NDF), $F$ denotes the Fresnel term, and $G$ denotes the geometric attenuation, $\mathbf{a}$ is the albedo for each Gaussian, $m$ denotes the metallics for each Gaussian. We can therefore separate the whole $\mathbf{L}_{o}(\mathbf{w}_{o})$ into two separate parts (specular light $\mathbf{L}_{s}$ in equation. 10 and diffuse light $\mathbf{L}_{d}$ in equation. 11). We utilize the split-sum approximation [101] to approximate the two terms separately. By approximation, we can separate the integral of the light and BRDF product into two separate integrals, which we can pre-integrate separately for rapid querying.
For the specular light term $\mathbf{L}_{s}$, we have
[…]
The first term denotes the integral of the specular BRDFs. It only depends on the $\mathbf{w}_{i}\cdot\mathbf{n}$ and the roughness $r$ and is independent of the illumination. The second term denotes the integral of ingoing radiance with Specular NDF $D$, and it depends on the reflection direction $\mathbf{r}=2(\mathbf{w}_{i}\cdot\mathbf{n})\mathbf{n}-\mathbf{w}_{i}$ and roughness $r$. It can also be pre-integrated and queried from a filtered cubemap.
For the diffuse light term $\mathbf{L}_{d}$, we have:
[…]
The integration term can also be efficiently calculated with pre-integration and queried from a cubemap.
Occlusion Calculation. As there exists a lot of self-occlusion during the animation of clothed human avatars, to achieve more realistic physically based rendering, we should consider the occlusion. However, calculating the occlusion with ray-tracing [51] for each Gaussian is time-consuming and hinders rendering speed. To achieve occlusion calculation while also maintaining the fast rendering nature of Gaussian Splatting, we proposed to approximate the occlusion as the ambient occlusion from SMPL [39] mesh $M(\boldsymbol{\theta},\boldsymbol{\beta})=(\mathbf{V},\mathbf{F})$. Baking the ambient occlusion of the SMPL mesh can reduce the time for occlusion calculation. Specifically, we first query the nearest surface points $\mathbf{v}_{k}^{*}$ from mesh $M(\boldsymbol{\theta},\boldsymbol{\beta})$ for each 2D Gaussian $\mathbf{p}_{k}^{*}$ at articulated space. We then utilize the surface points $\mathbf{v}_{k}^{*}$ and its corresponding surface normal $\mathbf{n}_{k}^{*}$ from mesh to bake the ambient occlusion for the $k$-th Gaussian:
[…]
where the $V_{\mathbf{v}_{k}^{*},\mathbf{w}}\in[0,1]$ is the visibility function at each surface points $\mathbf{v}_{k}^{*}$. The integral is approximated by casting rays in $100$ random directions around each vertex. The visibility of the specular light is, therefore, $1-AO_{k}$.
The overall outgoing radiance for each Gaussian is:
$\mathbf{c}_{k}^{\text{PBR}}=\mathbf{L}_{o}(\mathbf{w}_{o})=\mathbf{L}_{d}+(1-AO_{k})\mathbf{L}_{s}$ (13)
here the $\mathbf{w}_{o}$ denotes the outgoing direction, corresponding to the view direction for each 2D gaussian in world space.
Environment Map Representation. We utilize a $6\times 64\times 64$ trainable parameters as the environment map for Image-based lighting. We query the light at different mip-level. The base level signifies the pre-integrated lighting for the minimum supported roughness value. Subsequent smaller mip-levels are derived from this base level using pre-filtering approach [101].
Figure 3: Visualization of our PBR optimization stage. At the initial stage, both the D-normal $\mathbf{N}$ (Normal from depth points) and the R-normal (Rendered Splat Normal) are not accurate. After the Geometry Fix Stage ($\lambda_{\text{NC}_{2}}=0,\lambda_{\text{NC}_{1}}=1$). The Geometry is repaired while the splat normal is still messy. We then set ( $\lambda_{\text{NC}_{1}}=0,\lambda_{\text{NC}_{2}}=1$), and finally achieve that both the geometry and splat normal are consistent and accurate.
Physical based Rendering with splatting. We first project the 2D Gaussians into image space with the camera parameters. We can then calculate the PBR color $\mathbf{c}_{k}^{PBR}$ for each 2D Gaussian and apply the splatting as
$\mathbf{c}^{\text{PBR}}(x)=\sum_{i=1}\mathbf{c}_{i}^{\text{PBR}}o_{i}\mathcal{G}_{i}(\mathbf{u}(x))\prod_{j=1}^{i-1}(1-o_{j}\mathcal{G}_{j}(\mathbf{u}(x)))$ (14)
we can also get the corresponding albedo, roughness, metallic, and visibility maps by following the same splatting process of 2DGS.
3.4 Reconstruct animatable PBR-aware 2DGS from monocular videos
Our training process consists of two stages. In the first stage, we apply an Image Reconstruction loss to capture the rough shape of the dynamic humans. In the second stage, we use Physically-Based Rendering (PBR) to estimate PBR materials and achieve accurate surface reconstruction. We add additional regularization terms to obtain a suitable solution for such a highly ambiguous problem. Specifically, given $M$ frames from a monocular video and the estimated poses of the human $\{(\mathbf{I}_{1},\boldsymbol{\theta}_{1}),\dots,(\mathbf{I}_{M},\boldsymbol{\theta}_{M})\}$ we optimize $\mathbf{G},\widehat{\mathbf{W}},\widetilde{\mathbf{B}},\widetilde{\mathbf{W}}$ with the ground-truth images.
Training Stage I – Image Reconstruction. In the first stage, we aim to reconstruct the rough shape for clothed avatars with view-dependent color represented as spherical harmonics. We apply the image reconstruction loss, regularization term on these Gaussian parameters, and the normal-depth consistency loss proposed by [33]. Denote the rendered image as $\hat{\mathbf{I}}$, the training loss at this stage is:
$L_{\text{RF}}=\ell_{1}(\hat{\mathbf{I}},\mathbf{I})+\lambda_{\text{SSIM}}L_{\text{SSIM}}(\hat{\mathbf{I}},\mathbf{I})+L_{\text{reg}}+\lambda_{\text{NC}}L_{\text{NC}}$ (15)
we only render the foreground avatar and apply the $L_{\text{RF}}$ loss on the foreground avatar and do not model the background. Due to Gaussians’ discrete representation nature, without any regularization term, the reconstructed cloth avatars may be overfitting to the sparse 2D observation. We therefore follow previous works [40] and regularize the standard variation $\sigma_{k}$ of Gaussian attributes in the KNN neighborhood of each Gaussian $\mathbf{G}_{k}$, which leads to:
$L^{\sigma}_{k}=\sum_{\text{attr}_{k}\in\{\mathbf{R}_{k},\mathbf{S}_{k},o_{k},\mathbf{c}_{k},\Delta\mathbf{W}_{k}\}}\lambda_{\text{attr}_{k}}\sigma(\text{attr}_{k})$ (16)
where $\sigma_{\text{attr}_{k}}$ denotes the standrd variation of the Gaussian attributes in the KNN neighborhood. Furthermore, we encourage learnable blending weight, and the latent bones for cloth animation for each Gaussian should also be small, thus alleviating the overfitting for the cloth deformation. We follow the previous work [35] to store the learnable blending weight in a 3D voxel and apply a tri-linear interpolation to query the blending weight $\Delta W_{k}$ for each Gaussian. We therefore regularize the learnable blending weight to be small for both the SMPL bones and latent bones with:
$L_{k}^{\text{norm}}=\lambda_{\widehat{\mathbf{W}}}\|\Delta\mathbf{W}_{k}\|_{2}+\lambda_{\widetilde{\mathbf{W}}}\|\widetilde{\mathbf{W}}_{k}\|_{2}$ (17)
The total regularization loss is:
$L_{\text{reg}}=\frac{1}{N_{\text{gs}}}\sum_{k=1}^{N_{\text{gs}}}L_{k}^{\sigma}+L_{k}^{\text{norm}}$ (18)
To ensure the orientation of each 2D Gaussians locally aligned with the actual surface and suitable for physically based rendering. We follow 2DGS [33] and align the splat’s normal with the gradients of the depth maps as follows:
$L_{\text{NC}}=\sum_{i}w_{i}(1-\mathbf{n}_{i}^{T}\mathbf{N})$ (19)
where $w_{i}=o_{i}\mathcal{G}_{i}(\mathbf{u}(x))\prod_{j=1}^{i-1}(1-o_{j}\mathcal{G}_{j}(\mathbf{u}(x)))$ denotes the blending weight of the $i$-th intersection, $\mathbf{n}_{i}$ is the normal of the splat facing the camera, $\mathbf{N}$ is normal derived from depth map, In contrast to the $\ell_{1}$ loss proposed by previous works [47, 49], we find this normal-consistency loss can achieve more smooth surface and consistent orientation for each gaussian along the ray, as validated in our ablation study.
Training Stage II – PBR optimization. At the PBR stage, we first apply the image reconstruction loss on the PBR image. Denote the rendered image of PBR as $\mathbf{\hat{I}}_{\text{pbr}}(\mathbf{a},m,r,\mathbf{n})$, which is computed from albedo, materials, and splat normal. We apply:
$L_{\text{PBR}}=\ell_{1}(\hat{\mathbf{I}}_{\text{pbr}},\mathbf{I})+\lambda_{\mathbf{a}}L_{\text{smooth}}(\mathbf{a})+$ (20)
$\lambda_{r}L_{\text{smooth}}(r)+\lambda_{m}L_{\text{smooth}}(m)+$
$\lambda_{\text{NC}}L_{\text{NC}^{*}}+\lambda_{\text{white}}L_{\text{white}}$
Here the $L_{\text{NC}^{*}}$ is slightly different from the $L_{\text{NC}}$ used in the previous stage. With original normal-consistency loss from the previous stage, we can not reconstruct totally accurate geometry, especially under the extreme high light conditions as shown in Fig 3 and original 2DGS [33] paper. At the PBR stage, the appearance of Gaussians are correlated with the splat normal. Thus we can further improve the geometry during the optimization. Therefore, we can not freeze the Gaussian shape during the PBR optimization stage as proposed by GS-IR [47]. However, if we directly optimize with Equation 19, the splat normal still needs to be consistent with the inaccurate geometry and can not improve the geometry of clothed avatars. If we discard this normal consistency constraint, the splat normal will be totally messy during the optimization. Therefore, we propose a progressive optimization strategy to achieve both accurate geometry and consistent splat normal with following:
$L_{\text{NC}^{*}}=\lambda_{\text{NC}_{1}}\sum_{i}w_{i}(1-\text{sg}(\mathbf{n}_{i}^{T})\mathbf{N})$ (21)
$+\lambda_{\text{NC}_{2}}\sum_{i}w_{i}(1-\mathbf{n}_{i}^{T}\mathbf{N})$
where $\text{sg}(\cdot)$ denotes the stop gradient. At the initial stage, we set $\lambda_{\text{NC}_{2}}=0,\lambda_{\text{NC}_{1}}=1$, which means we utilize the splat normal to optimize the geometry $\mathbf{N}$. Then we set $\lambda_{\text{NC}_{1}}=0,\lambda_{\text{NC}_{2}}=1$ to force the splat normal to be aligned with the fixed geometry. We visualize the splat normal and depth derived normal during the optimization stage at Fig 3. To overcome the ambiguity between the lighting and PBR materials and obtain a plausible solution for PBR materials, we follow  [48] to utilize a smoothness loss which forces the materials and albedo to be smooth at the planar surfaces where the image gradients are absent. This can be written as:
$L_{\text{smooth}}(\mathbf{a})=\frac{1}{N_{\text{pixel}}}\sum_{i,j}|\partial_{\mathbf{x}}\mathbf{a}_{ij}|e^{-\|\partial_{\mathbf{x}}\mathbf{I}_{ij}\|}$ (22)
$+\frac{1}{N_{\text{pixel}}}\sum_{i,j}|\partial_{\mathbf{y}}\mathbf{a}_{ij}|e^{-\|\partial_{\mathbf{y}}\mathbf{I}_{ij}\|}$
where $N_{\text{pixel}}$ denotes the numbe of pixels for the image. I represents the ground-truth images, $\partial_{\mathbf{x}},\partial_{\mathbf{y}}$ represents the gradients of an image or material map in $\mathbf{x}$ and $\mathbf{y}$ directions. We apply the smoothness loss on the rendered albedo, metallic and roughness map. Furthermore, We assume a near-natural white incident light to obtain a plausible lighting materials disentangle. We therefore apply the white light regularizations on the environment map as:
$L_{\text{white}}=\sum_{\text{channel}}(\mathbf{L}_{\text{channel}}-\frac{1}{3}\sum_{\text{channel}}\mathbf{L}_{\text{channel}}),{\text{channel}}\in\{R,G,B\}$ (23)
3.5 Novel pose animation under novel illuminations
After optimization, we can get the canonical PBR-aware 2DGS $\mathbf{G}$. To achieve novel pose animation under novel illuminations, we first deform the optimized canonical PBR-aware 2DGS to the observation space with the given novel pose which is unseen in the training process following the equation 7. We then use the environment map of the novel illuminations to build a cubemap. For each viewpoint, we query the cubemap to calculate the PBR color for each Gaussian at observation space with the equation 13. Finally, we apply the splatting process (equation 14) to render the avatar at given novel illuminations.
4 Experiments
4.1 Evaluation datasets
To validate the effectiveness of our proposed methods, we use a synthetic dataset (RANA [92]) and two real-world datasets (PeopleSnapshot [104], ZJU-MoCap [38, 105]) for performance evaluation.
RANA [92]. RANA [92] is a synthetic dataset that includes the ground-truth albedo, normal, and relighting results at novel pose for evaluation. The training view is a subject with an A-pose rotating in front of a camera in the scene. The testing set includes the same subject under unknown illumination with random poses. To quantitatively evaluate the physical material of the reconstructed human, we use 8 subjects from the RANA dataset [92] by following the same protocols in IntrinsicAvatar [29].
PeopleSnapshot [104] PeopleSnapshot captures subjects that hold a nearly A-pose and rotate in front of the camera. The subjects are under natural illumination from the real world. We follow previous works [40, 35] and select 4 subjects from the dataset for qualitative evaluation. After optimization, our reconstructed avatar can be rendered in a novel pose under novel illuminations.
ZJU-MoCap [38, 105] In ZJU-MoCap, the subjects captured in the video may conduct very complex motions. The shadow effect during the complex motion challenges our physically based rendering. We use the dataset to show that our method can achieve human intrinsic disentangles from a monocular video with complex human motions.
4.2 Baselines
We compare our method with two recently proposed methods, IntrinsicAvatar [29] and R4D [25]. IntrinsicAvatar [29] is the most recent state-of-the-art method with publicly available training codes for the physically based inversed rendering of clothed humans. We also choose the Relighting4D (R4D) as our baselines [25]. Since the original R4D implementation does not utilize the alpha loss, we follow [29] to report a variant of R4D* [^2] The training code of RelightableAvatar [30] are not fully released with guidance at the current stage, which we encountered some errors when run the code. They only release one checkpoint on their MobileStage dataset for inference. Thus we do not compare our method with it. that utilize an alpha loss.
4.3 Evaluation Metrics
Figure 4: Qualitative Results on the RANA Dataset. We visualize the PBR image, albedo, and normal from the training view. We then relight the clothed human avatars at 2 different novel poses under the novel illuminations.
TABLE I: Quantitative comparison to the baselines on the RANA dataset. The bold font indicates the best performance. The underline indicates the second-best performance. Our method costs minimal training time.
  Method Relighting (novel pose) Normal Error $\downarrow$ Albedo Training Time
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
R4D [25] 16.41 0.8023 0.2213 42.69 18.24 0.7780 0.2414 2 days
R4D* [25] 18.32 0.8132 0.1913 27.38 18.23 0.8254 0.2043 2 days
IntrinsicAvatar [29] 20.68 0.8818 0.1139 9.98 22.23 0.8862 0.1615 4 hours
Ours 21.30 0.8871 0.1134 10.33 21.76 0.8908 0.1492 40 mins
Figure 5: Qualitative comparisons on the PeopleSnapshot dataset. We visualize the PBR image, albedo, and normal from the training view. We then relight the clothed human avatars at training pose and novel pose under the novel illumination.
We use the following metrics for performance evaluation on the synthetic dataset RANA [92].
Albedo: PSNR/SSIM/LPIPS: We use the image quality metrics to evaluate the albedo estimation from the training view. Since ambiguity exists between the albedo scale and light intensity, we follow previous works [22, 29] to align the estimated albedo with a scale of the ground-truth albedo.
Normal Estimation Error: The metric evaluates the normal errors between the predicted normal and the ground-truth normal from the training view. We empirically find that the rendered normal is better.
Relighting (novel pose): PSNR/SSIM/LPIPS: We render the clothed human in the novel pose under the novel illumination from the test set. We then evaluate the image quality between the predicted images and the ground-truth.
We only compared our method with baselines qualitatively on the real-world datasets since there is no ground-truth albedo and material information.
4.4 Experimental Results
Figure 6: Qualitative results on the ZJU-MoCap [38] dataset. In the left 6 columns, we show the rendered image, estimated albedo, roughness, metallic, normal, and the approximated occlusion. In the right 3 columns, we visualize the relighting results under different novel illuminations at different poses.
Figure 7: Our clothed human avatars in novel views under novel illuminations. The avatars are from the PeopleSnapshot [104] dataset. We visualize the avatars from different views with different poses and under different unknown illuminations from [106].
Training details. We first train 5,000 steps for the image reconstruction stage. We apply the normal-depth consistency regularizer after 3,000 steps, from which we can start to align the splat normal to the optimized geometry. We then conduct 3,000 steps to optimize the PBR materials. The overall training step is 8,000 steps, which consumes about 40 minutes on a single Nvidia 3090Ti GPU (with peak 13GB GPU memory usage), which is almost 6 times faster than IntrinsicAvatar. We set the $\lambda_{\text{NC}}=0.05$ for the normal-consistency loss. The weights in the smooth loss ($\lambda_{\mathbf{a}},\lambda_{r},\lambda_{m}$) are set to be 0.02. The weight of $\lambda_{\text{white}}$ is set to be 0.1. We follow the learning rate in the original Gaussian attribute in Gaussian Splatting [32, 33]. For the additional parameters of the PBR attribute and the learnable environment light, we set the learning rate the same as that of the spherical harmonics appearance. In the image reconstruction stage, the overall loss is eq. 15, while in the PBR optimization stage, the overall loss is eq. 20.
Rendering time. The rendering time mainly depends on the number of Gaussians. The time cost mainly comes from the occlusion calculation. For PeopleSnapshot [104] with $540\times 540$ resolution, there are $\sim 50,000$ Gaussians after optimization, the physically based rendering time is 0.2 seconds per image, which is faster than all previous neural-rendering based works (20s per image as shown in the Fig. 4).
4.4.1 Comparison on synthetic dataset
We compare with the baselines in Tab.I 333 We re-run the baselines [25, 29] with the official implementation provided by the authors:
https://github.com/FrozenBurning/Relighting4D,
https://github.com/taconite/IntrinsicAvatar on our dataset splits and report the reproduced both the quantitative and qualitative results. Our reported relighting results are higher than those reported in previous works as we employ a similar processing method to albedo by dilating the foreground outward by 100 pixels.
and Fig.4. We can observe that R4D* [25] cannot reconstruct the surface geometry accurately. Since such a method relies on the time-conditioned embedding during optimization, thus, it cannot generalize to the novel poses. Moreover, its implicit representation makes it lack of the details. While Intrinsic Avatar [29] can generalize to the novel pose and its multi-level Hash-Grids [31] representation provide much more details, the albedo details (in the man’s face on the left of Fig. 5) are still missing. For the female on the right of Fig. 5, the dress surface reconstructed by  [29] tends to overfit the texture, while our methods can reconstruct the smooth surface for the dress. Furthermore, our methods can bring high light during relighting, which brings superiority for relighting (novel pose). The quantitative results in Tab. I also demonstrate this. Compared with R4D and R4D*, our methods reduce $75.8\%$ and $62.6\%$ normal errors and significantly improve the PSNR of estimated albedo by $19.2\%$. Compared with IntrinsicAvatar, our methods achieve comparable estimated albedo quality (a lower PSNR (0.46 dB), but a higher SSIM and LPIPS) and more accurate relighting quality (a higher PSNR (0.62 dB)). Further, compared with SOTA methods, our methods can achieve a 5×5$\times$5 × faster in training speed and a 100×100$\times$100 × faster in rendering speed.
4.4.2 Comparison on real-world dataset
We also compare our method with the baselines on the real-world nearly A-pose dataset qualitatively in Fig. 5. The novel poses are selected from the AIST dataset [107]. We can observe that the albedo estimation of R4D* [25] fails to recover the plausible solution, its geometry is inaccurate. The relighting results at both the training and test poses are unsuitable for the environmental light. Compared with Intrinsic Avatar [29], our methods can produce more plausible relighting results in addition to the 5×5$\times$5 × faster in training speed and 100×100$\times$100 × faster in rendering speed.
4.4.3 PBR properties estimation results on the real-world datasets with complex motions
We follow the setting in the InstantNVR [34] and choose view 4 from ZJU-MoCap as the training view. We visualize the intrinsic estimation results in Fig. 6. Our methods can produce plausible BRDF estimation results. We also render the clothed human at novel illumination and get well-reproduced results. We also visualize the approximated occlusion calculated from the template mesh to show the effectiveness of our proposed human avatar occlusion approximation strategy. We can observe from Fig. 6 that our method can achieve a plausible PBR properties estimation from a monocular video with complex human motions, and can also re-render them with an appearance suitable to the environment lighting.
4.4.4 Reposed clothed human avatars in novel views under novel illuminations
We select the motion sequence from the CAPE dataset [108, 95] and repose our reconstructed clothed human avatars. We then select 3 different views and re-render each avatar under 3 different environment lighting conditions. The results in Fig. 7 demonstrate the robustness of our methods. The reconstructed avatars from our methods can generalize well to the out-of-distribution pose under totally novel illuminations while maintaining 3D consistency. We provide a video in our supplementary to provide a temporally consistent visualization.
Figure 8: Ablation studies of different training strategies. We visualize the Albedo estimation, D-Normal (normal from depth points), and the R-Normal (rendered normal of each splat).
4.5 Ablation studies
We conduct ablation studies on different representations (3DGS or 2DGS), different normal consistency regularization($\ell_{1}$-loss or cosine similarity loss), the effectiveness of occlusion calculation, the effectiveness of the occlusion approximation strategies, the PBR regularization loss, and our progressive training strategy.
4.5.1 Training strategy
TABLE II: Ablation study of our proposed training strategy
  Method Albedo Normal Error $\downarrow$
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Freeze shape 18.78 0.7661 0.2581 21.51
w/o prog. train 20.21 0.8247 0.1977 12.89
Ours SGIA 23.17 0.8839 0.1701 11.31
We first compare different training strategies for the PBR optimization stage in the Tab. II and the Fig. 8. ”Freeze shape” means that we choose the optimization strategy of GS-IR [47], which freeze the shape after image reconstruction stage and only optimize the PBR materials of each Gaussian. ”w/o progressive” means that we do not utilize the strategy in (equation 21) and directly utilize the normal-consistency loss (equation 19) from the first stage. We can observe from the Fig. 8 that if we utilize the ”freeze shape”, the geometry of Gaussian is fixed. The geometry from the first stage cannot bring any improvement. If we do not utilize our proposed progressive training strategy, the normal of each Gaussian is always enforced to align with the inaccurate geometry in the first stage. The inaccurate geometry will hurt the geometry optimization in the PBR stage. With our proposed progressive training strategy, both the normal of each splat and the geometry of the avatar can be aligned consistently to the accurate geometry. Moreover, the inaccurate geometry also leads to inaccurate albedo estimation. The quantitative results in Tab. II further demonstrate these.
4.5.2 Primitive representation
Figure 9: Ablation studies of different primitive choices. We visualize the rendering image, albedo, D-Normal (normal from depth points), and the R-Normal (rendered normal of each splat).
TABLE III: Ablation study of different primitive representation
  Method Albedo Normal Error $\downarrow$
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
PBR-aware 3DGS 22.07 0.8739 0.1586 12.33
PBR-aware 2DGS 22.43 0.8923 0.1545 11.71
We ablate different choices of primitive representation in Tab. III and Fig. 9. Since original 3DGS [32] do not have an attribute for splat orientation, we follow the previous relightable 3DGS [47, 50] to assign a learnable normal vector for each Gaussian and keep all the other strategy unchanged. We can observe from Fig. 9 that with 3DGS as a primitive representation, though the splat normal (R-normal) can be very smooth, there exists a lot of uneven region in the Depth (D-normal). The R-normal and the D-Normal are not consistent. However, with the 2DGS representation, the R-Normal and the D-Normal can be both consistent and accurate. The Normal Error in Tab. III also demonstrates this quantitatively.
4.5.3 Normal-consistency loss
TABLE IV: Ablation study of different consistency loss
  Method Albedo Normal Error $\downarrow$ Relighting (novel pose)
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
$\ell_{1}$-loss 21.76 0.8798 0.1746 20.33 15.87 0.7844 0.2162
cosine-loss 22.43 0.8923 0.1545 11.71 19.77 0.8642 0.1317
Figure 10: Ablation studies of different normal-consistency loss choices. We visualize the albedo, the D-Normal (normal from depth points), the R-Normal (rendered normal of each splat), and the relighting results with novel poses under a novel illumination.
We ablate the normal consistency loss in Tab IV and Fig. 10. ”$\ell_{1}$-loss” means that we follow the previous relightable 3DGS [50, 47] and utilize the $\ell_{1}$-loss to minimize the difference of the linear blended normal and the normal from depth points. We keep the remaining factors unchanged. We can observe that utilizing the $\ell_{1}$-loss, the normal of each splat cannot align with the D-Normal well, while the inaccurate splat normal also leads to bad relighting results as shown in the Fig. 10. The quantitative results from Tab IV also provide evidence.
Figure 11: Ablation study of our proposed occlusion calculation. We visualize the Albedo and Relighting results under novel illuminations, both at novel poses. The red rectangle indicates the coupled occlusion information in albedo. The green rectangle indicates the decoupled occlusion information from the albedo.
TABLE V: Ablation study of the effectiveness of the occlusion calculation
  Method Albedo Normal Error $\downarrow$
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
w/o occlusion 21.08 0.8616 0.1711 11.87
w/ occlusion 21.31 0.8701 0.1702 10.69
4.5.4 Occlusion calculation
We ablate our proposed occlusion calculation in Tab. V and Fig. 11. We can observe that the shadow information may be mixed in the albedo without the occlusion calculation during the optimization stage. With the occlusion calculation, our methods successfully disentangle the occlusion information from the albedo. The quantitative results in Tab. V also prove the effectiveness of decoupling the shadow effect and the PBR properties of our occlusion approximation strategy.
4.5.5 Occlusion calculation from template mesh or clothed mesh
We ablate our approximation of the occlusion calculation from the template mesh in Tab. VI and Fig. 12. For the fully clothed mesh case, we follow 2DGS [33] and use the TSDF Fusion [109] to extract the mesh of the clothed human avatar at the canonical pose. We query the points by using tri-linear interpolation from the canonical skinning weight voxel to get the skinning weight for each canonical vertex. We keep other regularization terms the same as those in 2DGS and optimize the same steps.Tab. VI shows that the accuracy of the estimated PBR properties with the occlusion calculated from the template mesh only has a minor gap to the occlusion calculated from the fully clothed mesh. However, the rendering time cost using the fully clothed mesh is about 10×10$\times$10 × longer than that with the template mesh. This is because the number of vertices and faces in the fully clothed mesh can be 20×20$\times$20 × more than those in the template mesh. Fig. 12 shows that the calculated visibility map and the estimated albedo do not have any significant difference. This demonstrates that our proposed occlusion approximation strategy can achieve a faster rendering speed without sacrificing the accuracy of the estimated PBR properties.
Figure 12: Ablation studies of different occlusion calculation strategies. At the top of the figure, we visualize different canonical meshes used in different occlusion calculation strategies. At the bottom of the figure, we visualize the albedo and the visibility map of different occlusion calculation strategies both at novel poses.
TABLE VI: Ablation study of the occlusion calculation from the template mesh or the clothed mesh
  Method Albedo Rendering Time $\downarrow$
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Full clothed mesh 22.01 0.8831 0.1267 2s
Template mesh (Ours) 21.93 0.8827 0.1259 0.2s
4.5.6 PBR regularization
We ablate our PBR regularization choices in the Tab. VII and Fig. 13. ”w/o. white reg” means that we do not utilize the white light regularization in Equation 23. ”w/o. smooth reg” means that we do not utilize the smooth regularization in Equation 22. We can observe from Fig. 13 that without the light regularization, the albedo estimation and the splat normal can be noisy. Without the smooth regularization, the albedo estimation does not include details (such as eyes on the face), the splat normal and the roughness are totally messy. The roughness value varied even in the region that should have the same material properties. The PBR rendering image may also fit concisely with the ground-truth image. The quantitative results in Tab. VII also demonstrate the effectiveness of regularization terms.
Figure 13: Ablation study of PBR Regularization. We visualize the rendering image, albedo, R-Normal (rendered normal of each splat), and the roughness.
TABLE VII: Ablation study of PBR regularization
  Method Albedo Rendering Time $\downarrow$
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
w/o white reg 22.05 0.9045 0.1761 10.71
w/o smooth reg 21.83 0.9046 0.1863 10.93
Ours Full 22.44 0.8874 0.1758 9.4
5 Conclusion And Future Work
In this paper, we propose SGIA, a relightable clothed human avatar representation that can achieve fast reconstruction of the PBR properties from monocular videos. In contrast to all previous NeRF-based works that rely on Monte-Carlo sampling to integrate the in-going radiance, our methods can be easily equipped with pre-integration to achieve efficient lighting computation. We approximate the occlusion by ray-casting from template mesh to balance accuracy and speed. Furthermore, to achieve accurate geometry reconstruction guided by PBR, we propose a simple but effective progressive optimization strategy that improves the geometry and then aligns the splat normal with the geometry. Overall, our methods can achieve five times the training speed and 100 times the rendering speed compared with previous SOTA approaches while maintaining high-quality PBR properties estimation and realistic re-lighting results.
While our method efficiently reconstructs the PBR properties of clothed human avatars from monocular videos, it is important to acknowledge its limitations. One significant limitation is the lack of detailed facial expression capture. Our approach does not leverage expression capture techniques, resulting in less detailed facial animations for the avatars. Future work could enhance realism and expressiveness by integrating facial expression techniques, as explored in studies such as [110, 111, 112], with our full-body gesture reconstruction. A promising direction is to combine the expressive Gaussian Avatar representation with our full PBR representation pipeline to achieve both relightable and expressive human avatar reconstruction.
Acknowledgments
The authors would like to thank Umar Iqbal and Shaofei Wang for providing details about preprocessing RANA dataset, Shaofei Wang for providing the scripts of the camera trajectory setting for free viewpoint rendering on CAPE dataset, Zibo Zhao and Shenhan Qian for the valuable discussion and feedbacks at the initial stage of the project, Yahao Shi and Yanmin Wu for providing the implementation details of GIR, Jingwei Xu for providing the guidance of the draft pipeline figure.